OpenStudy (anonymous):

Which method is correct when finding the product of (x + 2)2? Method 1 (x + 2)^2 x2 + 2^2 x^2 + 4 Method 2: (x + 2)2 (x + 2)(x + 2) x^2 + 4x + 4

OpenStudy (anonymous):

method 2

OpenStudy (campbell_st):

the 2nd is correct as the process of squaring something means multiply by itself e.g 5^2 = 5 x 5 so (x+2)^2 = (x +2)(x+2)

OpenStudy (anonymous):

yeah method 2